a2+b2=25 and a3+b3=91 find the value of a and b.
Answer by Alan3354(69443)

a2+b2=25 and a3+b3=91 find the value of a and b.
----------
Use ^ (Shift 6) for exponents.
---
a^2+b^2=25 and a^3+b^3=91 find the value of a and b.
---
3 & 4